# DMVFC: Deep Learning Based Functionally Consistent Tractography Fiber Clustering Using Multimodal Diffusion MRI and Functional MRI

### v1

@@ I Introduction
I-A Tractography Parcellation Using Fiber Clustering
@@
I-B Functional and Microstructural Information Along WM Fiber Tracts
@@
I-C Unsupervised Deep Feature Learning and Multi-view Clustering
@@ I-D Contributions
II METHODS
The overall workflow of DMVFC is shown in Figure 1. There are two main training stages in DMVFC: (1) multi-view pretraining and (2) collaborative fine-tuning. In the first stage, the multi-view pretraining module computes two kinds of embeddings from fiber geometric information and brain functional signals (Section II-A). In the second stage, the collaborative fine-tuning module refines the pretrained embeddings to ensure that the clustering outcomes integrate both dMRI and fMRI information simultaneously (Section II-B). During the inference stage (Section II-C), fractional anisotropy (FA) is incorporated as supplementary information to enhance the clustering process.
II-A Multi-view Pre-training for Embedding Computation
@@ II-A1 Network inputs
For capturing functional signals, fibers are represented as point clouds defined by the BOLD signals at the endpoints. The fiber endpoints are located in the grey matter (except for tracts such as the CST, where one endpoint is within the white matter brainstem region).To improve cross-subject comparability and measurement robustness, we use the BOLD signals at the two endpoints near the cortex as the functional representation: endpoints lie near the cortex, provide a higher signal-to-noise ratio, and facilitate parcellation-based alignment. By contrast, correlations of along-tract white-matter BOLD can be susceptible to local confounds—e.g., small vessel/venous signals, crossing-fiber interference and partial-volume effects, field inhomogeneity, and regional variability in the hemodynamic response function—leading to larger phase/latency variability and noise and thus hampering cross-subject consistency assessments. Each endpoint contains 1,200 time points (see Section III), which were randomly downsampled to 600 to enable more efficient CPU processing. As a result, for a fiber $i$, its functional signal input is formed as $\mathbf{x}_{i}^{2}\in\mathbb{R}^{2\times 600}$.
II-A2 Network training
@@
II-B Collaborative Fine-tuning
@@
It is notable that the initialization of the centroids of clusters at the very beginning of fine-tuning process in both geometric information view and functional information view are crucial for the final results of optimization of the model. We calculate the centroids of geometric information embeddings first, and then use the fiber indexes of centroids in geometric information view as the centroids of the functional information embedding. The impact of the initialization strategy is further examined in the ablation study presented in Section III-C.
@@
II-C Inference Stage
During the inference stage, for a new subject, we first compute the FA similarity between individual fibers. In parallel, the geometric input (as introduced in Section II-A1) is constructed and fed into the network to obtain the corresponding embeddings. The Manhattan distance is employed to quantify the FA similarity between fibers. This similarity value is then scaled by a factor of 30 and added to the geometric distance produced by the model to compute the final distance. This final distance is subsequently used to estimate the probability of each input belonging to a given cluster, as described in Eq. 5 (Section II-B). The final cluster assignment for each fiber is determined by selecting the cluster with the highest probability. Notably, fMRI data are not needed at the inference stage. The model operating in the geometric information view is trained with guidance from functional data, but performs prediction independently of it.
III EXPERIMENTS AND RESULTS
III-A Data Acquisition and Preprocessing
@@
TABLE I: Average GM correlation and $\alpha$ in different bundles under different model.
@@
III-B Implementation
III-C Experimental Results
III-C1 Model performance comparison with state-of-the-art methods
To evaluate our clustering results, we compare our method with three state-of-the-art fiber clustering methods: QuickBundles (QB) [3], a Riemannian Framework for Structurally Curated Functional Clustering of Brain WM Fibers [41], and Deep Fiber Clustering (DFC) [13]. In brief, QB is an efficient fiber clustering technique that groups similar fibers based on extracted features and spatial proximity. The Riemannian Framework is an advanced method for functional fiber clustering using WM BOLD signals along fibers. DFC is a deep learning-based method on which our method is built (see Section II-A).
@@ III-C1 Model performance comparison with state-of-the-art methods
III-C2 Ablation study of embedding initialization before fine-tuning
As discussed in Section II-B, the initialization of the clustering embeddings plays a critical role in the subsequent fine-tuning process. To systematically evaluate its impact, we perform an ablation study designed to analyze the influence of different initialization strategies on the final model’s performance. This investigation aims to provide empirical insights into how initialization affects model representation quality. Our method utilizes centroids of geometric information view to initialize the centroids of functional information view. In the ablation study, individual initialization method without explicit initialization relies on performing initialization within their respective embedding spaces by using centroids derived from each individual view. All other parameters and experimental conditions are held constant to isolate the effect of initialization strategy on model performance. Following this initialization approach, both model configurations undergo fine-tuning for 20 epochs. The metrics fMRI endpoints correlation and $\alpha$ are calculated on the test datasets. This design enables a clear assessment of how omitting the specific initialization impacts the fine-tuning process and final outcomes. As shown in Table LABEL:table_initial, our method achieves higher fMRI correlation and less $\alpha$. It indicates that the method of utilizing geometric information embedding centroids to initialize functional information view has a positive effect on subsequent fine-tuning.
@@ III-C2 Ablation study of embedding initialization before fine-tuning
III-C3 Ablation study of incorporation of fMRI and FA information
To evaluate the impact of incorporatin fMRI and FA information on model performance, we conducted a comparative analysis involving three distinct model configurations relative to a baseline model: (1) the proposed model incorporating both FA and fMRI data, (2) the baseline model enhanced solely with fMRI data, and (3) the baseline model augmented exclusively with FA information. The fMRI-enhanced model was implemented using the same multi-view collaborative clustering framework as the proposed model, with the FA component removed to isolate the effect of fMRI data. Conversely, the FA-only model is based on the DFC [13] framework, incorporating FA information exclusively during the inference stage, as detailed in Section II-C. These results are summarized in Figure 3, where DFC[13] serves as the reference baseline, and improvements in fMRI correlation are depicted via bar plots. The findings clearly indicate that integrating FA information leads to consistent and significant enhancements in model performance across all configurations compared to the baseline. This suggests that FA provides valuable complementary information, thereby improving the robustness and accuracy of brain functional modeling. Moreover, the proposed model integrating both fMRI and FA information demonstrates superior improvement compared to models incorporating each modality individually. Notably, in certain fiber bundles, the performance improvement exceeds the additive effect of combining the separate improvements observed from fMRI and FA alone, indicating a synergistic interaction between the two data types.
TABLE II: Ablation study of centroid initialization in the embedding space. fMRI correlation and alpha metrics for the final results are reported for models using different initialization strategies before fine-tuning.
@@ III-C3 Ablation study of incorporation of fMRI and FA information
TABLE III: Hausdorff Distance between corresponding functional pathways, fibers in clusters, and fibers in bundles (MD denotes Mean Hausdorff Distance)
@@
III-C4 Experiments on clustering consistency across subjects
@@
III-C5 Visualization of fiber clustering results
@@
III-C6 Visualization of embedding space

### v2

@@ 1 Introduction
1.1 Tractography Parcellation Using Fiber Clustering
@@
1.2 Functional and Microstructural Information Along WM Fiber Tracts
@@
1.3 Unsupervised Deep Feature Learning and Multi-view Clustering
@@ 1.4 Contributions
2 METHODS
The overall workflow of DMVFC is shown in Figure 1. There are two main training stages in DMVFC: (1) multi-view pretraining and (2) collaborative fine-tuning. In the first stage, the multi-view pretraining module computes two kinds of embeddings from fiber geometric information and brain functional signals (Section 2.1). In the second stage, the collaborative fine-tuning module refines the pretrained embeddings to ensure that the clustering outcomes integrate both dMRI and fMRI information simultaneously (Section 2.2). During the inference stage (Section 2.3), fractional anisotropy (FA) is incorporated as supplementary information to enhance the clustering process.
2.1 Multi-view Pre-training for Embedding Computation
@@ 2.1.1 Network inputs
For capturing functional signals, fibers are represented as point clouds defined by the BOLD signals at the endpoints. The fiber endpoints are located in the grey matter (except for tracts such as the CST, where one endpoint is within the white matter brainstem region).To improve cross-subject comparability and measurement robustness, we use the BOLD signals at the two endpoints near the cortex as the functional representation: endpoints lie near the cortex, provide a higher signal-to-noise ratio, and facilitate parcellation-based alignment. By contrast, correlations of along-tract white-matter BOLD can be susceptible to local confounds—e.g., small vessel/venous signals, crossing-fiber interference and partial-volume effects, field inhomogeneity, and regional variability in the hemodynamic response function—leading to larger phase/latency variability and noise and thus hampering cross-subject consistency assessments. Each endpoint contains 1,200 time points (see Section 3), which were randomly downsampled to 600 to enable more efficient CPU processing. As a result, for a fiber $i$, its functional signal input is formed as $\mathbf{x}_{i}^{2}\in\mathbb{R}^{2\times 600}$.
2.1.2 Network training
@@
2.2 Collaborative Fine-tuning
@@
It is notable that the initialization of the centroids of clusters at the very beginning of fine-tuning process in both geometric information view and functional information view are crucial for the final results of optimization of the model. We calculate the centroids of geometric information embeddings first, and then use the fiber indexes of centroids in geometric information view as the centroids of the functional information embedding. The impact of the initialization strategy is further examined in the ablation study presented in Section 3.3.
@@
2.3 Inference Stage
During the inference stage, for a new subject, we first compute the FA similarity between individual fibers. In parallel, the geometric input (as introduced in Section 2.1.1) is constructed and fed into the network to obtain the corresponding embeddings. The Manhattan distance is employed to quantify the FA similarity between fibers. This similarity value is then scaled by a factor of 30 and added to the geometric distance produced by the model to compute the final distance. This final distance is subsequently used to estimate the probability of each input belonging to a given cluster, as described in Eq. 5 (Section 2.2). The final cluster assignment for each fiber is determined by selecting the cluster with the highest probability. Notably, fMRI data are not needed at the inference stage. The model operating in the geometric information view is trained with guidance from functional data, but performs prediction independently of it.
3 EXPERIMENTS AND RESULTS
3.1 Data Acquisition and Preprocessing
@@
3.2 Implementation
Table 1: Average GM correlation and $\alpha$ in different bundles under different model.
@@ 3.2 Implementation
3.3 Experimental Results
3.3.1 Model performance comparison with state-of-the-art methods
To evaluate our clustering results, we compare our method with three state-of-the-art fiber clustering methods: QuickBundles (QB) [3], a Riemannian Framework for Structurally Curated Functional Clustering of Brain WM Fibers [5], and Deep Fiber Clustering (DFC) [14]. In brief, QB is an efficient fiber clustering technique that groups similar fibers based on extracted features and spatial proximity. The Riemannian Framework is an advanced method for functional fiber clustering using WM BOLD signals along fibers. DFC is a deep learning-based method on which our method is built (see Section 2.1).
@@ 3.3.1 Model performance comparison with state-of-the-art methods
3.3.2 Ablation study of embedding initialization before fine-tuning
As discussed in Section 2.2, the initialization of the clustering embeddings plays a critical role in the subsequent fine-tuning process. To systematically evaluate its impact, we perform an ablation study designed to analyze the influence of different initialization strategies on the final model’s performance. This investigation aims to provide empirical insights into how initialization affects model representation quality. Our method utilizes centroids of geometric information view to initialize the centroids of functional information view. In the ablation study, individual initialization method without explicit initialization relies on performing initialization within their respective embedding spaces by using centroids derived from each individual view. All other parameters and experimental conditions are held constant to isolate the effect of initialization strategy on model performance. Following this initialization approach, both model configurations undergo fine-tuning for 20 epochs. The metrics fMRI endpoints correlation and $\alpha$ are calculated on the test datasets. This design enables a clear assessment of how omitting the specific initialization impacts the fine-tuning process and final outcomes. As shown in Table LABEL:table_initial, our method achieves higher fMRI correlation and less $\alpha$. It indicates that the method of utilizing geometric information embedding centroids to initialize functional information view has a positive effect on subsequent fine-tuning.
3.3.3 Ablation study of incorporation of fMRI and FA information
To evaluate the impact of incorporatin fMRI and FA information on model performance, we conducted a comparative analysis involving three distinct model configurations relative to a baseline model: (1) the proposed model incorporating both FA and fMRI data, (2) the baseline model enhanced solely with fMRI data, and (3) the baseline model augmented exclusively with FA information. The fMRI-enhanced model was implemented using the same multi-view collaborative clustering framework as the proposed model, with the FA component removed to isolate the effect of fMRI data. Conversely, the FA-only model is based on the DFC [14] framework, incorporating FA information exclusively during the inference stage, as detailed in Section 2.3. These results are summarized in Figure 3, where DFC[14] serves as the reference baseline, and improvements in fMRI correlation are depicted via bar plots. The findings clearly indicate that integrating FA information leads to consistent and significant enhancements in model performance across all configurations compared to the baseline. This suggests that FA provides valuable complementary information, thereby improving the robustness and accuracy of brain functional modeling. Moreover, the proposed model integrating both fMRI and FA information demonstrates superior improvement compared to models incorporating each modality individually. Notably, in certain fiber bundles, the performance improvement exceeds the additive effect of combining the separate improvements observed from fMRI and FA alone, indicating a synergistic interaction between the two data types.
Table 2: Ablation study of centroid initialization in the embedding space. fMRI correlation and alpha metrics for the final results are reported for models using different initialization strategies before fine-tuning.
@@ 3.3.3 Ablation study of incorporation of fMRI and FA information
Table 3: Hausdorff Distance between corresponding functional pathways, fibers in clusters, and fibers in bundles (MD denotes Mean Hausdorff Distance)
@@
3.3.4 Experiments on clustering consistency across subjects
@@
3.3.5 Visualization of fiber clustering results
@@
3.3.6 Visualization of embedding space
@@ 5 CONCLUSION
Acknowledgements
This work is in part supported by the National Key R&D Program of China (No. 2023YFE0118600), and the National Natural Science Foundation of China (No. 62371107).